OpenStudy (anonymous):

fill in the blank sin a ____ =1/2 (sin(a+b)+sin(a-b)) options: A)sin a B)sin b C)cos b D)cos a

OpenStudy (anonymous):

please help! :)

OpenStudy (anonymous):

cos b

OpenStudy (anonymous):

sin(a+b)+sin(a-b)=(sin a*cos b+sin b*cos a)+(sin a*cos b-sin b*cos a)=2*sin a*cos b

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can you give me a gold medal?